Project

General

Profile

Bug #24593

HWR Resonance control cables

Added by Brian Chase 4 months ago. Updated 3 months ago.

Status:
Resolved
Priority:
Normal
Start date:
07/02/2020
Due date:
% Done:

100%

Estimated time:
Spent time:
Duration:

Description

There is still confusion about cables.

From Joe's notes.

Powered up cavity #5 at low power, and began calibration process, locked onto cavity, field < 1 MV/m. However the tuner did not function (see Phil's entry # 174633), though the cavity frequency was close enough so that adjusting he LO frequency allowed the system to lock on. Investigation of the tuner issue was not fruitful, so we decided to change to cavity #4. Saw the same behavior with cavity 4 - able to power up (low power, 20W), and lock to the cavity, establish field. But tuner did not function (no change in frequency observed). Also we didn't see the pressure transducer output voltage change. So more investigation is needed to understand what is going on. We were able to successfully operate the tuner for cavity #3 last week, Also all tuners were checked out and run over their entire ranges (> 100kHz) at 2K.

A few of us will meet tomorrow AM to discus the path forward. It may be likely that we move ahead with cavity characterization in SEL mode (with minor adjustment to LO frequency if required) so that we can establish cavity performance.

History

#1 Updated by Brian Chase 4 months ago

At least cables 5 and 6 are labeled wrong. Resonance control is also on the wrong labview page.

#2 Updated by Shrividhyaa Sankar Raman 4 months ago

  • % Done changed from 0 to 80
  • Tuner pod cables labeling have been checked for cavities 3 through 8.
  • Cable database needs to be updated with these changes.

#3 Updated by Shrividhyaa Sankar Raman 4 months ago

  • Status changed from New to Work in progress

#4 Updated by Brian Chase 3 months ago

There are four pairs of cables that can be in one of two states. When can we say that we have the cables correctly labeled and connected correctly?

#5 Updated by Philip Varghese 3 months ago

  • % Done changed from 80 to 100
  • Status changed from Work in progress to Resolved

The confusion about tuner cable is definitively resolved when both the tuner is operated an RF applied
to the cavity. At this point all cavities have been subjected to this combined test and the final confusion
was corrected last week when it was observed that cavity 7 and 8 were swapped. The cables were relabled after the
test.



Also available in: Atom PDF